# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 4.1 BPS locus
We parametrize Abelian gauge or flavour fields on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ as follows:
@@
with ${\mathfrak{m}}\in\mathbb{Z}$ labelling the flux through the spindle base and $h=0,\dots,{\left({n-1}\right)}$ the $\mathbb{Z}_{n}$ holonomy along the Hopf fiber. We again gauged away the component ${\mathcal{A}}_{x}{\left({x}\right)}$. As in the case of $\mathbb{\Sigma}\times S^{1}$, the behaviour of ${\mathcal{A}}_{\varphi}$ is
@@
with $\mathfrak{m}_{\pm}=\mathfrak{m}\,\mathfrak{a}_{\pm}$ and ${n_{+}}{\mathfrak{a}_{-}}-{n_{-}}{\mathfrak{a}_{+}}=1$.
If we do not impose reality conditions on fields, the vector-multiplet BPS equations $\delta\lambda=\delta\widetilde{\lambda}=0$ require that on the supersymmetric locus the dynamical scalar $\sigma$ and the auxiliary scalar $D$ are
$\sigma|_{\rm BPS}=\sigma_{c}+\frac{{\rm i}}{b_{1}b_{2}}{\left[{-{n_{+}}b_{1}h+{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}{\mathcal{A}}_{\varphi}{\left({x}\right)}}\right]}~{},$
@@ 4.2 One-loop determinants
We calculate the one-loop determinant for a chiral multiplet on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ via the same technique adopted in the case of $\mathbb{\Sigma}\times S^{1}$. Especially, the eigenfunction $B_{m_{\varphi},m_{\psi}}{\left({x}\right)}$ is non-singular at $x=0,\pi/2$ if the integers ${\left({m_{\varphi},m_{\psi}}\right)}$ do not violate the following bounds:
${n_{+}}m_{\varphi}\leq{\mathfrak{m}_{+}}~{},$ ${n_{-}}{\left({m_{\varphi}-nm_{\psi}+h}\right)}\leq{\mathfrak{m}_{-}}~{}.$ (4.5)
@@
$m_{\varphi}$ $=-n\,\ell_{\varphi}-{\left\llbracket{h+{\left\llbracket{\Delta_{+}}\right\rrbracket_{n}}+j_{\psi}}\right\rrbracket_{n}}+{\left\lfloor{\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}~{},$
$m_{\psi}$ $=\ell_{\psi}-\ell_{\varphi}+{\left\lfloor{\frac{\Delta_{+}}{n}}\right\rfloor}+{\left\lfloor{\frac{h+{\left\llbracket{\Delta_{+}}\right\rrbracket_{n}}+j_{\psi}}{n}}\right\rfloor}~{},$
$\Delta_{\pm}$ =±$={\left\lfloor{\pm\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}-{\left\lfloor{\pm\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}~{},\qquad\ell_{\psi},\ell_{\varphi}\in\mathbb{N}~{},\qquad j_{\psi}=0,\dots,{\left({n-1}\right)}~{}.$ (4.6)
Correspondingly, the eigenfunction $\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}$ is non-singular at $x=0,\pi/2$ if the integer inequalities
${n_{+}}n_{\varphi}\geq{\mathfrak{m}_{+}}~{},$ ${n_{-}}{\left({n_{\varphi}-nn_{\psi}+h}\right)}\geq{\mathfrak{m}_{-}}~{},$ (4.7)
hold. They are fulfilled if
$n_{\varphi}$ $=n\,k_{\varphi}+t_{\varphi}-{\left\lfloor{-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}~{},$
$n_{\psi}$ $=-k_{\psi}+k_{\varphi}+{\left\lfloor{-\frac{\Delta_{-}}{n}}\right\rfloor}+{\left\lfloor{\frac{h+{\left\llbracket{-\Delta_{-}}\right\rrbracket_{n}}+t_{\varphi}}{n}}\right\rfloor}~{},$
$\ell_{\psi},\ell_{\varphi}\in\mathbb{N}~{},\qquad t_{\varphi}=0,\dots,{\left({n-1}\right)}~{}.$ (4.8)
@@
$Z^{\mathcal{B},\Phi}_{\text{1-L}}$ $=\prod_{j=0}^{n-1}\prod_{k,\ell\in\mathbb{N}}\frac{\frac{1}{b}{\left({k-\frac{1}{n}{\left\lfloor{\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}+\frac{{\left\llbracket{h+{\left\llbracket{\Delta_{+}}\right\rrbracket_{n}}+j}\right\rrbracket_{n}}}{n}}\right)}+b{\left({\ell-\frac{1}{n}{\left\lfloor{\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}+\frac{j}{n}}\right)}+\frac{Q}{2n}+\frac{{\rm i}\,u}{n}}{\frac{1}{b}{\left({k-\frac{1}{n}{\left\lfloor{-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}+\frac{j}{n}}\right)}+b{\left({\ell-\frac{1}{n}{\left\lfloor{-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}+\frac{{\left\llbracket{h+{\left\llbracket{-\Delta_{-}}\right\rrbracket_{n}}+j}\right\rrbracket_{n}}}{n}}\right)}+\frac{Q}{2n}-\frac{{\rm i}\,u}{n}}~{},$
$b$ $=\sqrt{\frac{{n_{-}}b_{1}}{{n_{+}}b_{2}}}~{},\qquad Q={\left({\frac{b^{-1}}{{n_{+}}}+\frac{b}{{n_{-}}}}\right)}~{},\qquad u=\sqrt{\frac{b_{1}b_{2}}{{n_{+}}{n_{-}}}}\,\sigma_{c}-{\rm i}{\left({1-r}\right)}\frac{Q}{2}-{\rm i}\,b\,h~{}.$ (4.9)
Hence, on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ the one-loop determinant for a chiral and a vector multiplet with weights $\rho\in\mathfrak{R}_{G}$ and $\alpha\in{\rm adj}\,G$ for a gauge group $G$ read
$Z^{\rm CM}_{\text{1-L}}=\prod_{\rho\in\mathfrak{R}_{G}}Z^{\mathcal{B},\Phi}_{\text{1-L}}{\left({\rho{\left({\mathfrak{m}_{\pm}}\right)},\rho{\left({u}\right)},\rho{\left({h}\right)}}\right)}~{},$ $Z^{\rm VM}_{\text{1-L}}=\prod_{\alpha\in{\rm adj}\,G}Z^{\mathcal{B},\Phi}_{\text{1-L}}{\left({\alpha{\left({\mathfrak{m}_{\pm}}\right)},\alpha{\left({u}\right)},\alpha{\left({h}\right)}}\right)}|_{r=2}~{}.$ (4.10)
These match the one-loop determinants on the squashed lens space computed in [31] upon setting $n_{\pm}=1$ and employing the function
@@
Eventually, if $n_{\pm}=n=1$, the one-loop determinant above reduces to that of the squashed three-sphere. In this limit the one-loop determinant above also matches the determinant of fluctuations for matter fields computed in [32] on the resolved space of the branched sphere, once the identifications $b_{1}=q_{\rm there}$ and $b_{2}=p_{\rm there}$ are made.
5 Spindle from the squashed three-sphere
@@
In this section we derive one-loop determinants for matter and gauge fields on the pure two-dimensional spindle starting from three dimensions. Indeed, by exploiting the fact that the spindle non-trivially embeds into the squashed three-sphere $S^{3}_{q,p}$, we reduce the one-loop determinant associated with an $\mathcal{N}=2$ chiral multiplet on $S^{3}_{q,p}$ to the one-loop determinant of an $\mathcal{N}=(2,2)$ chiral multiplet on $\mathbb{\Sigma}$. In turn, this will provide the one-loop determinant of a $\mathcal{N}=(2,2)$ vector multiplet.
5.1 $SL(2,\mathbb{Z})$ reduction
5.1.1 General Construction
Let the change of coordinates from $\theta\in{\left[{0,\pi/2}\right]}$ and $\phi_{1,2}\in{\left[{0,2\pi}\right]}$ to coordinates $\Theta\in{\left[{0,\pi}\right]}$ and $\sigma,y\in{\left[{0,2\pi}\right]}$ be
$\theta$ $=\Theta/2~{},$
$\phi_{1}$ $=t_{11}\,\sigma+t_{12}\,y~{},$
$\phi_{2}$ $=t_{21}\,\sigma+t_{22}\,y~{},$ (5.1)
with $t\in SL(2,\mathbb{Z})$. The inverse change of coordinates reads
$\Theta$ $=2\theta~{},$
$\sigma$ $=t_{22}\phi_{1}-t_{12}\phi_{2}~{},$
$y$ $=-t_{21}\phi_{1}+t_{11}\phi_{2}~{}.$ (5.2)
Also,
$\partial_{\theta}$ $=2\partial_{\Theta}~{},$
$\partial_{\phi_{1}}$ $=t_{22}\partial_{\sigma}-t_{21}\partial_{y}~{},$
$\partial_{\phi_{2}}$ $=-a_{12}\partial_{\sigma}+a_{11}\partial_{y}~{},$ (5.3)

### v2

@@ 4.1 BPS locus
We parametrize Abelian gauge or flavour fields on ${\mathbb{\Lambda}}$ as follows:
@@
with ${\mathfrak{m}}\in\mathbb{Z}$ labelling the flux through the spindle base and $h=0,\dots,{\left({n-1}\right)}$ the $\mathbb{Z}_{n}$ holonomy along the Hopf fiber. As we do not impose reality conditions on the BPS locus, supersymmetry allows for $\mathfrak{f}_{G}\neq 0$ and $\mathfrak{m}\neq 0$, in principle. Then, the BPS locus (4.1) includes and generalizes those employed e.g. in [43, 32], which are recovered in the subcase $\mathfrak{m}=0$. In fact, we will explicitly check that the integer deformation tuned by $\mathfrak{m}$ is perfectly under control and has the effect of shifting the value of the holonomy $h$ in the final result of the partition function. Similar shifts were observed in [49], where vortex configurations on the Hopf base of $S^{3}$ were considered. Notice that we again gauged away the component ${\mathcal{A}}_{x}{\left({x}\right)}$. As in the case of $\mathbb{\Sigma}\times S^{1}$, the behaviour of ${\mathcal{A}}_{\varphi}$ is
@@
with $\mathfrak{m}_{\pm}=\mathfrak{m}\,\mathfrak{a}_{\pm}$ and ${n_{+}}{\mathfrak{a}_{-}}-{n_{-}}{\mathfrak{a}_{+}}=1$. In this context, the configurations characterized by ${\left({{\mathfrak{m}_{+}},{\mathfrak{m}_{-}}}\right)}$ assume a role akin to that of $H$ in [49].
If we do not impose reality conditions on fields, then the locus (4.1) is supersymmetric with respect to the vector-multiplet BPS equations $\delta\lambda=\delta\widetilde{\lambda}=0$, which require that on the supersymmetric locus the dynamical scalar $\sigma$ and the auxiliary scalar $D$ are
$\sigma|_{\rm BPS}=\sigma_{c}+{\rm i}{\left({{n_{+}}b_{1}^{-1}+{n_{-}}b_{2}^{-1}}\right)}{\mathcal{A}}_{\varphi}{\left({x}\right)}~{},$
@@ 4.2 One-loop determinants
We calculate the one-loop determinant for a chiral multiplet on ${\mathbb{\Lambda}}$ via the same technique adopted in the case of $\mathbb{\Sigma}\times S^{1}$. Especially, the eigenfunction $B_{m_{\varphi},m_{\psi}}{\left({x}\right)}$ is non-singular at $x=0,\pi/2$ if the integers ${\left({m_{\varphi},m_{\psi}}\right)}$ do not violate the following bounds:
$-{n_{+}}m_{\varphi}+n\,{t_{+}}m_{\psi}-q_{G}{t_{+}}h+q_{G}{\mathfrak{m}_{+}}\geq 0~{},$ ${n_{-}}m_{\varphi}-n\,{t_{-}}m_{\psi}+q_{G}{t_{-}}h-q_{G}{\mathfrak{m}_{-}}\leq 0~{}.$ (4.5)
@@
$m_{\varphi}$ $={t_{-}}{\left({q_{G}{\mathfrak{m}_{+}}-n\ell_{\varphi}-j_{\varphi}}\right)}+{t_{+}}{\left({n\ell_{\psi}+j_{\psi}-q_{G}{\mathfrak{m}_{-}}}\right)}~{},$
$m_{\psi}$ $={n_{+}}\ell_{\psi}-{n_{-}}\ell_{\varphi}+{\left\lfloor{\frac{{n_{+}}j_{\psi}+q_{G}{\left({h-\mathfrak{m}}\right)}}{n}}\right\rfloor}-{\left\lfloor{\frac{{n_{-}}j_{\varphi}}{n}}\right\rfloor}~{},$
${\left\llbracket{{n_{-}}j_{\varphi}}\right\rrbracket_{n}}$ $={\left\llbracket{{n_{+}}j_{\psi}+q_{G}{\left({h-\mathfrak{m}}\right)}}\right\rrbracket_{n}}~{},$
$\ell_{\psi},\ell_{\varphi}$ $\in\mathbb{N}~{},\qquad j_{\varphi},j_{\psi}=0,\dots,{\left({n-1}\right)}~{}.$ (4.6)
Correspondingly, the eigenfunction $\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}$ is non-singular at $x=0,\pi/2$ if
$n_{\varphi}$ $={t_{-}}{\left({n\ell_{\varphi}+k_{\varphi}+q_{G}{\mathfrak{m}_{+}}}\right)}-{t_{+}}{\left({n\ell_{\psi}+k_{\psi}+q_{G}{\mathfrak{m}_{-}}}\right)}~{},$
$n_{\psi}$ $={n_{-}}\ell_{\varphi}-{n_{+}}\ell_{\psi}+{\left\lfloor{\frac{{n_{-}}k_{\varphi}+q_{G}{\left({h-\mathfrak{m}}\right)}}{n}}\right\rfloor}-{\left\lfloor{\frac{{n_{+}}k_{\psi}}{n}}\right\rfloor}~{},$
${\left\llbracket{{n_{+}}k_{\psi}}\right\rrbracket_{n}}$ $={\left\llbracket{{n_{-}}k_{\varphi}+q_{G}{\left({h-\mathfrak{m}}\right)}}\right\rrbracket_{n}}~{},$
$k_{\varphi},k_{\psi}$ $=0,\dots,{\left({n-1}\right)}~{}.$ (4.7)
@@
$Z^{\mathcal{B},\Phi}_{\text{1-L}}$ $=\prod_{j_{1}=0}^{n-1}\prod_{k,\ell\in\mathbb{N}}\frac{\prod_{j_{2}\in\mathbb{J}_{-}{\left({j_{1},\mathfrak{h}}\right)}}{\left[{\frac{1}{b}{\left({k+\frac{j_{2}}{n}}\right)}+b{\left({\ell+\frac{j_{1}}{n}}\right)}+\frac{Q}{2n}+\frac{{\rm i}u}{n}}\right]}}{\prod_{j_{3}\in\mathbb{J}_{+}{\left({j_{1},\mathfrak{h}}\right)}}{\left[{\frac{1}{b}{\left({k+\frac{j_{1}}{n}}\right)}+b{\left({\ell+\frac{j_{3}}{n}}\right)}+\frac{Q}{2n}-\frac{{\rm i}u}{n}}\right]}}~{},$
$b$ $=\sqrt{\frac{b_{1}}{b_{2}}}~{},\qquad Q=\frac{1}{b}+b~{},\qquad\mathfrak{h}=h-\mathfrak{m}~{},$
$u$ $=q_{G}\sqrt{b_{1}b_{2}}{\left[{\sigma_{c}+{\rm i}{\left({\frac{{\mathfrak{m}_{+}}}{b_{1}}+\frac{{\mathfrak{m}_{-}}}{b_{2}}}\right)}}\right]}-{\rm i}{\left({1-r}\right)}\frac{Q}{2}~{},$
$\mathbb{J}_{\pm}{\left({j_{1},\mathfrak{h}}\right)}$ =$={\left\{{j_{0}=0,\dots,{\left({n-1}\right)}:{\left\llbracket{n_{\pm}j_{0}}\right\rrbracket_{n}}={\left\llbracket{n_{\mp}j_{1}+q_{G}\mathfrak{h}}\right\rrbracket_{n}}}\right\}}~{}.$ (4.8)
Hence, on ${\mathbb{\Lambda}}$ the one-loop determinant for a chiral and a vector multiplet with weights $\rho\in\mathfrak{R}_{G}$ and $\alpha\in{\rm adj}\,G$ for a gauge group $G$ read
$Z^{\rm CM}_{\text{1-L}}$ $=\prod_{\rho\in\mathfrak{R}_{G}}Z^{\mathcal{B},\Phi}_{\text{1-L}}{\left({\alpha{\left({u}\right)},\alpha{\left({\mathfrak{h}}\right)}}\right)}~{},$
$Z^{\rm VM}_{\text{1-L}}$ $=\prod_{\alpha\in{\rm adj}\,G}Z^{\mathcal{B},\Phi}_{\text{1-L}}{\left({\alpha{\left({u}\right)},\alpha{\left({\mathfrak{h}}\right)}}\right)}|_{r=2}~{}.$ (4.9)
These match the one-loop determinants on the squashed lens space computed in [32] upon setting $n_{\pm}=1$, $\mathfrak{h}=h_{\rm there}$ and employing the function
@@
Further setting $b=1$ makes our computation agree with the analysis conducted in [43] for the unsquashed lens space. Eventually, if $n_{\pm}=n=1$, the one-loop determinant above reduces to that of the squashed three-sphere. In this limit the one-loop determinant above also matches the determinant of fluctuations for matter fields computed in [33] on the resolved space of the branched sphere, once the identifications $b_{1}=q_{\rm there}$ and $b_{2}=p_{\rm there}$ are made.
5 Spindles from the squashed three-sphere
@@
In this section we derive one-loop determinants for matter and gauge fields on the pure two-dimensional spindle starting from three dimensions. Indeed, by exploiting the fact that the spindle non-trivially embeds into a squashed three-sphere $S^{3}_{b_{s},b_{c}}$, we reduce the one-loop determinant associated with an $\mathcal{N}=2$ chiral multiplet on $S^{3}_{b_{s},b_{c}}$ to the one-loop determinant of an $\mathcal{N}=(2,2)$ chiral multiplet on $\mathbb{\Sigma}$. We first focus on chiral multiplets coupled to $U(1)$ vector multiplets. Eventually, we shall consider one-loop determinants of vector multiplets, corresponding to chiral multiplets in the representation $\mathfrak{R}_{G}={\rm adj}\,G$ with $R$-charge $r=2$. The method that we employ below is different from analogous reduction techniques discussed in [34] and extended in [55, 56].
5.1 General construction
Let
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+b_{s}^{2}\sin^{2}x\,{\rm d}\varphi_{1}^{2}+b_{c}^{2}\cos^{2}x\,{\rm d}\varphi_{2}^{2}~{},$ (5.1)
be the line element on the squashed three-sphere $S^{3}_{b_{s},b_{c}}$, with $f$ being the squashing function
$f=-\sqrt{b_{s}^{2}\cos^{2}x+b_{c}^{2}\sin^{2}x}~{},$ (5.2)
while the ranges of the coordinates are $x\in{\left[{0,\pi/2}\right]}$ and $\varphi_{1},\varphi_{2}\in{\left[{0,2\pi}\right]}$. We can embed $S^{3}_{b_{s},b_{c}}$ in $\mathbb{C}^{2}$ by means of the complex coordinates
$z_{1}=b_{s}e^{{\rm i}\varphi_{1}}\sin x~{},\qquad z_{2}=b_{c}e^{{\rm i}\varphi_{2}}\cos x~{},$ $\frac{|z_{1}|^{2}}{b_{s}^{2}}+\frac{|z_{2}|^{2}}{b_{c}^{2}}=1~{}.$ (5.3)
We then define the following map from the coordinates ${\left({x,\varphi_{1},\varphi_{2}}\right)}$ to $\vartheta\in{\left[{0,\pi}\right]}$ and $\varphi,\psi\in{\left[{0,2\pi}\right]}$:
$x$ $={\left({\pi-\vartheta}\right)}/2~{},$
$\varphi_{1}$ $=n_{1}\psi+t_{1}\varphi~{},$
$\varphi_{2}$ $=n_{2}\psi+t_{2}\varphi~{},$ (5.4)
where ${\left({n_{1},n_{2},t_{1},t_{2}}\right)}\in\mathbb{Z}$, which satisfy $n_{1}t_{2}-n_{2}t_{1}=1$, can be interpreted as entries of a matrix in $SL(2,\mathbb{Z})$. The inverse change of coordinates reads
$\vartheta$ $=\pi-2\,x~{},$
$\varphi$ $=n_{1}\varphi_{2}-n_{2}\varphi_{1}~{},$
$\psi$ $=-t_{1}\varphi_{2}+t_{2}\varphi_{1}~{}.$ (5.5)